The freshman class at a college has a total of 250 students. One hundred and ten freshmen take math and 200 take English, while 75 take neither. How many freshman take both math and English?
Answer by ikleyn(52780)

.
You have an universal set of 250 elements and two its subsets, M (=Math) of 110 elements and E (=English) of 200 elements.


The union of the subsets M and E contains 250-75 = 175 elements.


Now use the formula


    n(M U E) = n(M) + n(E) - n(M n E)


for the numbers of elements.


From this formula,


    n(M n E) = n(M) + n(E) - n(M U E) = 110 + 200 - 175 = 135.


Answer.  135 freshman take both Math and English.

--------------

To see many other similar solved problems,  look into the lesson
    - Counting elements in sub-sets of a given finite set
in this site.